# DistPred: A Distribution-Free Probabilistic Inference Method for Regression and Forecasting

### v1

@@ 2.4 Incorporate alternative methodologies
DistPred is orthogonal to other methods, enabling its combination with alternative approaches to enhance estimation performance. Here, with a focus on computational efficiency and memory conservation, we opt to integrate MC Dropout with DistPred, thereby denoting the amalgamation as DistPred-MCD. In our experiments, we observed that DistPred-MCD can further enhance uncertainty quantification performance, albeit with a marginal increase in computational effort.
@@ 3.1 PICP and QICE metrics
In recent years, BNNs (Hernández-Lobato and Adams, 2015; Gal and Ghahramani, 2016) have emerged as a class of models that aims at estimating the uncertainty providing a more complete picture of $p(y|x)$. In (Han et al., 2022), it is contended that the metric utilized in BNNs to assess uncertainty estimates, namely the negative log-likelihood (NLL), is computed based on Gaussian densities. This assumption implies that they consider both the conditional distribution $p(y|x=x^{\prime})$ for all $x^{\prime}$ are Gaussian. However, this assumption is very difficult to verify for real-world datasets. We follow CARD (Han et al., 2022) and use the following two metrics, both of which are designed to empirically evaluate the degree of similarity between learned and true conditional distributions:
@@
$PICP:=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbbm{1}\{\hat{y}_{n}\leq q_{1-\alpha/2}\},$ (11)
where $q_{\alpha/2}$ and $q_{1-\alpha/2}$ represent the low and high percentiles, respectively, that we have selected for the predicted $\hat{y}$ outputs given the same $x$ input. This metric evaluates the proportion of accurate observations that lie within the percentile range of the generated $\hat{y}$ samples corresponding to each $x$ input. Within this study, we opt for the $2.5$th and $97.5$th percentiles, signifying that an optimal PICP value for the acquired model would ideally reach 95%.
However, a caveat of the PICP metric becomes apparent in the measurement of distribution differences. For instance, consider a situation where the 2.5th to 97.5th percentile range of the learned distribution encompasses the data falling between the 1st and 96th percentiles of the true distribution. Even with a sufficient number of samples, the resulting PICP value would still approach 95%. However, it becomes evident that a discrepancy exists between the learned distribution and the true distribution.
Drawing from this reasoning, Han et al. (2022) introduces a novel empirical metric called QICE. This metric can be perceived as an enhanced version of PICP, offering finer granularity and addressing the issue of uncovered quantile ranges. To calculate QICE, the initial step involves generating an adequate number of samples for each $\hat{y}$ value. These samples are then divided into $M$ bins of approximately equal sizes. Subsequently, the quantile values are determined at each boundary within these bins. The definition of QICE entails computing the mean absolute error (MAE) between the proportion of true data encompassed by each quantile interval and the optimal proportion, which is 1/M for all intervals:
$QICE:=\frac{1}{M}\sum_{m=1}^{M}|r_{m}-\frac{1}{M}|,\quad\text{where}\ \ r_{m}=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbbm{1}\{\hat{y}_{n}\leq q_{1-\alpha/2}\}.$ (12)
In this paper, we followed Han et al. (2022) and set $M=10$ for all experiments.
@@ 3.2 Toy examples
To demonstrate the effectiveness of DistPred, we initially conduct experiments on $8$ toy examples as done in CARD (Han et al., 2022). These examples are specifically crafted with distinct statistical characteristics in their data generating functions: some have a uni-modal symmetric distribution for their error term (linear regression, quadratic regression, sinusoidal regression), others have heteroscedasticity (log-log linear regression, log-log cubic regression) or multi-modality (inverse sinusoidal regression, 8 Gaussians, full circle).
The research demonstrates that a trained DistPred models has the capability to produce samples that closely resemble the true response variable for novel covariates. Additionally, it can quantitatively match the true distribution based on certain summary statistics. The study visualizes scatter plots comparing real and generated data for all eight tasks in Figure 5. In cases where the tasks involve unimodal conditional distributions, the interest region fills the region between the $2.5$th and $97.5$th percentiles of the generated $\hat{y}$ values.
@@ 3.3 UCI regression tasks
For experiments conducted on real-world datasets, we utilize the same $10$ UCI regression benchmark datasets (Asuncion and Newman, 2007) and follow the experimental protocol introduced by Hernández-Lobato and Adams (2015), which has also been followed by Gal and Ghahramani (2016) and Lakshminarayanan et al. (2017), as well as by Han et al. (2022). The dataset information can be found in Table 5 located in Appendix B.
@@
We compare DistPred with other state-of-the-art methods, including PBP (Hernández-Lobato and Adams, 2015), MC Dropout (Gal and Ghahramani, 2016), DeepEnsemble (Lakshminarayanan et al., 2017), and another deep generative model that estimates a conditional distribution sampler, GCDS (Zhou et al., 2021b), as well as diffusion model, CARD (Han et al., 2022). The multiple train-test splits are applied with a 90%/10% ratio, following the same methodology as Hernández-Lobato and Adams (2015) and Han et al. (2022) (20 folds for all datasets except 5 for Protein and 1 for Year). The reported metrics are presented as the mean and standard deviation across all splits. As pointed out by Han et al. (2022), we compare the QICE of different methods on various UCI datasets. Additional information regarding the experimental setup for these models is available in Appendix C. The experimental results, along with corresponding metrics, are presented in Table 1. The frequency with which each model achieves the best corresponding metric is reported in the penultimate row, while the frequency with which it achieves the top two positions is reported in the last row.
The results demonstrate that the DistPred method outperforms existing methods, often by a considerable margin. It is worth noting that these impressive results are achieved in a single forward pass of the DistPred method. Crucially, the performance of uncertain quantization can be enhanced even further by leveraging DistPred-MCD, a hybrid approach combining DistPred and MC Dropout.
The implementation of DistPred on UCI regression tasks follows a straightforward approach: We employ a basic multilayer perceptron (MLP) as the foundational framework, complemented by Equation 9 serving as the loss function for end-to-end training. Due to the fact that DistPred necessitates solely a single forward inference, its inference speed is notably rapid. Table 2 presents a comparison of the training and inference speeds of mainstream models. It should be noted that, for a fair comparison, the implementations of various models are constructed on the same backbone and utilized the same equipment. It is evident that DistPred is approximately $230$ times faster in training and about $90$ times faster in inference compared to the state-of-the-art model CARD. The inference speed of DistPred is slower than its training speed because it involves calculating distribution statistical metrics like QICE and PICP.
Table 2: Comparison of model training and inference times (minutes) on UCI boston datasets.
Models DistPred DistPred-MCD PBP MC Dropout CARD
Inference 0.095 $\pm$ 4E-3 0.027 $\pm$ 5E-3 5.23 $\pm$ 0.10 4.62 $\pm$ 0.06 8.31 $\pm$ 0.17
@@ 3.4 Ablation study of the number of samples and ensembles
We investigate the influence of the number of samples generated by DistPred, as well as the number of ensembles of DistPred-MCD, on their respective performances. As shown in Figure 6, with an increase in the number of output samples and ensembles, the model’s performance shows a gradual improvement, eventually reaching a point of saturation.
Figure 6: Ablation study of the number of samples in DistPred (a) and the number of ensembles in DistPred-MCD (b).
3.5 Time series distribution forecasting
Baselines: We employ recent 10 SOTA methods for comparisons, including iTransformer Liu et al. (2023), PatchTST Nie et al. (2022), SCINet Liu et al. (2022a), TimesNet Wu et al. (2022), DLinear Zeng et al. (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021a), LogTrans Li et al. (2019) and Reformer Kitaev et al. (2020). We use the same experimental setup as (Zhou et al., 2021a) and (Liu et al., 2022a) and follow the same experimental protocol as (Zhou et al., 2021a). Univariate results can be found in Appendix D.
Datasets and setting: The detailed information pertaining to the datasets can be located in Appendix B. The models used in the experiments are evaluated over a wide range of prediction lengths to compare performance on different future horizons: 96, 192, 336 and 720. The experimental settings are the same for both multivariate and univariate tasks. We use the average of the MSE and MAE ($\frac{MSE+MAE}{2}$) to evaluate the overall performance of the model. It is noteworthy that DistPred provides an ensemble $\hat{Y}$ of response variables. Consequently, we employ the mean value of $\hat{Y}$ as the point estimate at that moment.
Main results: The results for multivariate TS forecasting are outlined in Table 3, with the optimal results highlighted in bold and the second-best results emphasized with underlined. It can be found that, despite not utilizing MSE and MAE, DistPred achieves state-of-the-art performance across all datasets and prediction length configurations. iTransformer and PatchTST stand out as the latest models acknowledged for their exceptional average performance. Compared with them, the proposed DistPred demonstrates an average performance increase of 3.5% and 16.5%, respectively, achieving a substantial performance improvement. We provide metrics, e.g., CRPS, QICE, PICP, for comparison by the future research community.
@@ 3.6 Visualization of the predictive distribution
The predictive distribution of DistPred is visualized in Figure 7. It can be observed that DistPred gives an ensemble of predictions (only top 10 are presented in the left subplot). Given all predictive ensemble values, the model can estimate the distribution of the response variable. Therefore, we can calculate confidence intervals at different levels, as shwon in the right subplot of Figure 7.
Figure 7: Visualization of the prediction results and the confidence intervals.
@@ 4 Conclusion
In this paper, we propose a novel method called DistPred, which is a distribution-free probabilistic inference method for regression and forecasting tasks. We transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. We also propose a hybrid method called DistPred-MCD, which combines DistPred with MC Dropout to further improve the performance of uncertain quantization. Experimental results demonstrate that DistPred outperforms existing methods, often by a considerable margin. We also extend time series forecasting from point estimation to distribution prediction and achieve state-of-the-art performance on multivariate and univariate time series forecasting tasks. In the future, we plan to extend DistPred to other tasks, such as classification and reinforcement learning.
Appendix A Related work
In supervised learning contexts, the endeavor to characterize the conditional distribution $p(y|x)$ beyond merely the conditional mean $\mathop{\mathbb{E}}[y|x]$ via deep neural networks has been a focal point of existing research efforts. These endeavors primarily concentrate on quantifying predictive uncertainty, with several approaches having been proposed. Bayesian neural networks (BNNs) represent one such approach, aiming to capture such uncertainty by positing distributions over network parameters, thereby encapsulating the model’s plausibility given the available data (Blundell et al., 2015; Hernández-Lobato and Adams, 2015; Gal and Ghahramani, 2016; Kingma et al., 2015; Tomczak et al., 2021). Another avenue is represented by Kendall and Gal (2017), which not only addresses uncertainties in model parameters but also incorporates an additive noise term into the neural network output to encompass uncertainties in model outputs. In parallel, ensemble-based methodologies (Lakshminarayanan et al., 2017; Liu et al., 2022b) have emerged to address predictive uncertainty. These methods involve amalgamating multiple neural networks with stochastic outputs, thereby providing a comprehensive approach to uncertainty quantification. Furthermore, the neural processes’ family (Garnelo et al., 2018b, a; Kim et al., 2019; Gordon et al., 2020) has introduced a suite of models tailored to capturing predictive uncertainty in a manner that extends beyond the distribution of available data, particularly tailored for few-shot learning scenarios.
Appendix B Dataset
B.1 Commonly used TS datasets

### v2

@@ 2.4 Incorporate Alternative Methodologies
DistPred is orthogonal to other methods, enabling its combination with alternative approaches to enhance estimation performance. Here, with a focus on computational efficiency and memory conservation, we opt to integrate MC Dropout with DistPred, thereby denoting the amalgamation as DistPred+MCD. In our experiments, we observed that using DistPred+MCD can further enhance uncertainty quantification performance, albeit with a marginal increase in computational effort.
@@ 3.1 PICP and QICE Metrics
The metric utilized in BNNs to assess uncertainty estimates, namely the negative log-likelihood (NLL), is computed based on Gaussian prior. This assumption implies that they consider both the conditional distribution $p(y|x=x^{\prime})$ for all $x^{\prime}$ are Gaussian. However, this assumption is very difficult to verify for real-world datasets. We follow (Han et al., 2022) and use the following two metrics, both of which are designed to empirically evaluate the degree of similarity between learned and true conditional distributions:
@@
$PICP:=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbb{I}\{\hat{y}_{n}\leq q_{1-\alpha/2}\},$ (14)
where $q_{\alpha/2}$ and $q_{1-\alpha/2}$ represent the low and high percentiles, respectively, that we have selected for the predicted $\hat{y}$ outputs given the same $x$ input. This metric evaluates the proportion of accurate observations that lie within the percentile range of the generated $\hat{y}$ samples corresponding to each $x$ input. Within this study, we opt for the $2.5$th and $97.5$th percentiles, signifying that an optimal PICP value for the model should ideally reach 95%.
However, a caveat of the PICP metric becomes apparent in the measurement of distribution differences. Drawing from this reasoning, Han et al. (2022) introduces a novel empirical metric called QICE. This metric can be perceived as an enhanced version of PICP, offering finer granularity and addressing the issue of uncovered quantile ranges. To calculate QICE, the initial step involves generating an adequate number of samples for each $\hat{y}$ value. These samples are then divided into $M$ bins of approximately equal sizes. Subsequently, the quantile values are determined at each boundary within these bins. The definition of QICE entails computing the mean absolute error (MAE) between the proportion of true data encompassed by each quantile interval and the optimal proportion, which is 1/M for all intervals:
$QICE:=$ $\frac{1}{M}\sum_{m=1}^{M}|r_{m}-\frac{1}{M}|,$ (15)
$\text{where}\ \ r_{m}=$ $\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\{\hat{y}_{n}\geq q_{\alpha/2}\}\cdot\mathbb{I}\{\hat{y}_{n}\leq q_{1-\alpha/2}\}.$
In this paper, we followed (Han et al., 2022) and set $M=10$ for all experiments.
@@ 3.2 Toy Examples
To demonstrate the effectiveness of DistPred, we initially conducted experiments on $8$ toy examples as done in (Han et al., 2022). These examples are specifically crafted with distinct statistical characteristics in their data generating functions: some have a uni-modal symmetric distribution for their error term (linear regression, quadratic regression, sinusoidal regression), while others exhibit heteroscedasticity (log-log linear regression, log-log cubic regression) or multi-modality (inverse sinusoidal regression, $8$ Gaussians, full circle).
The research demonstrates that a trained DistPred model has the capability to produce samples that closely resemble the true response variable for novel covariates. Additionally, it can quantitatively match the true distribution based on certain summary statistics. The study visualizes scatter plots comparing real and generated data for all eight tasks in Fig. 4. In cases where the tasks involve unimodal conditional distributions, the interest region fills the region between the $2.5$th and $97.5$th percentiles of the generated $\hat{y}$ values.
@@
Figure 5: DistPred and DistPred+MCD on UCI datasets.
Figure 6: Ablation studies of the number of samples in DistPred (a) and the number of ensembles in DistPred+MCD (b).
@@ 3.3 UCI Regression Tasks
Table 1: Comparison of model training and inference times (minutes) on UCI regression datasets when $K=1000$.
Models DistPred DistPred+MCD PBP MC Dropout CARD
Inference 0.026 $\pm$ 4E-3 0.027 $\pm$ 5E-3 5.23 $\pm$ 0.10 4.62 $\pm$ 0.06 8.31 $\pm$ 0.17
Figure 7: Comparison DistPred (ours) with conformal prediction methods. All results are averaged across all prediction lengths.
For experiments conducted on real-world datasets, we utilize the same $10$ UCI regression benchmark datasets (Asuncion and Newman, 2007) and follow the experimental protocol introduced by Hernández-Lobato and Adams (2015), which has also been followed by Gal and Ghahramani (2016) and Lakshminarayanan et al. (2017), as well as by Han et al. (2022). The dataset information can be found in Table 7 located in Appendix A.
@@
We compare DistPred with other state-of-the-art methods, including PBP (Hernández-Lobato and Adams, 2015), MC Dropout (Gal and Ghahramani, 2016), DeepEnsemble (Lakshminarayanan et al., 2017), and another deep generative model that estimates a conditional distribution sampler, GCDS (Zhou et al., 2021b), as well as the diffusion model, CARD (Han et al., 2022). The multiple train-test splits are applied with a 90%/10% ratio, following the same methodology as Hernández-Lobato and Adams (2015) and Han et al. (2022) (20 folds for all datasets except 5 for Protein and 1 for Year). The reported metrics are presented as the mean and standard deviation across all splits. As pointed out by Han et al. (2022), we compare the QICE of different methods on various UCI datasets. Additional information regarding the experimental setup for these models is available in Appendix A.2. The experimental results, along with corresponding metrics, are presented in Table 2. The frequency with which each model achieves the best corresponding metric is reported in the penultimate row, while the frequency with which it achieves the top two positions is reported in the last row.
The results demonstrate that DistPred outperforms existing methods, often by a considerable margin. It is worth noting that these impressive results are achieved in a single forward pass of the DistPred method. Crucially, as shown in Fig. 5, the performance of uncertain quantization can be further enhanced by leveraging DistPred+MCD, a hybrid approach that combines DistPred and MC Dropout.
The implementation of DistPred on UCI regression tasks follows a straightforward approach: We employ a basic multilayer perceptron (MLP) as the foundational framework, complemented by Eq. 13 serving as the loss function for end-to-end training. Due to the fact that DistPred necessitates solely a single forward inference, its inference speed is notably rapid. Table 1 presents a comparison of the training and inference speeds of mainstream models. It should be noted that, for a fair comparison, the implementations of various models are constructed on the same backbone and utilize the same equipment. It is evident that DistPred is approximately achieves at least 180x faster inference speed compared to state-of-the-art models. The inference speed of DistPred is slower than its training speed because it involves calculating distribution statistical metrics like QICE and PICP.
@@ 3.4 Ablation Study of The Number of Samples and Ensembles
We investigate the influence of the number of samples generated by DistPred, as well as the number of ensembles of DistPred+MCD, on their respective performances. In Fig. 6(a), we increase the number of samples of DistPred from 10 to 4000 to observe the changes in its QICE. In Fig. 6(b), we increase the number of ensembles of DistPred+MCD from 1 (DistPred) to 1000 to observe the changes in its QICE. It can be found that with an increase in the number of output samples and ensembles, the model’s performance shows a gradual improvement.
@@
3.5 Time Series Distribution Forecasting
Baselines: We employ recent 10 SOTA methods for comparisons, including iTransformer Liu et al. (2023), PatchTST Nie et al. (2022), SCINet Liu et al. (2022a), TimesNet Wu et al. (2022), DLinear Zeng et al. (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021a), LogTrans Li et al. (2019) and Reformer Kitaev et al. (2020). DistPred employs the same network architecture as iTransformer. We use the same experimental setup as (Zhou et al., 2021a) and (Liu et al., 2022a) and follow the same experimental protocol as (Zhou et al., 2021a). Univariate results can be found in Appendix E.
Datasets and Setting: The detailed information pertaining to the datasets can be located in Appendix A. The models Liang et al. (2024) used in the experiments are evaluated over a wide range of prediction lengths to compare performance on different future horizons: 96, 192, 336, and 720. The experimental settings are the same for both multivariate and univariate tasks. We use the average of the MSE and MAE ($\frac{MSE+MAE}{2}$) to evaluate the overall performance of the model. It is noteworthy that DistPred provides an ensemble $\hat{Y}$ of response variable. Consequently, we employ the mean value of $\hat{Y}$ as the point estimate at that moment.
Main Results: The results for multivariate TS forecasting are outlined in Table 3, with the optimal results highlighted in bold and the second-best results emphasized with underlined. It can be found that, despite not utilizing MSE and MAE, DistPred achieves state-of-the-art performance across all datasets and prediction length configurations. iTransformer and PatchTST stand out as the latest models acknowledged for their exceptional average performance. Compared with them, the proposed DistPred demonstrates an average performance increase of 3.5% and 16.5%, respectively, achieving a substantial performance improvement. We provide metrics, e.g., CRPS, QICE, PICP, for comparison by the future research community. Note that for long time series, computing these metrics on the entire testset can be very time-consuming and may lead to out-of-memory issues. Therefore, we propose calculating these metrics for each batch and then averaging the results.
3.6 Comparison with Conformal Prediction
Since conformal prediction offers an alternative approach for distribution prediction, we compare DistPred with conformal prediction methods on the time series forecasting tasks, including CQR (Romano et al., 2019), EnbPI (Xu and Xie, 2021), SPCI (Xu and Xie, 2023). Since PICP is the primary metric used by these methods, we adhere to their convention in this work. As shown in Fig. 7, DistPred achieves better performance than conformal prediction methods on all datasets and prediction lengths, with an average PICP improvement of 4.5%. This further underscores the high competitiveness of DistPred.
Table 4: Stability experiments of loss function (Eq. 13).
1% 25% 50% 75% 100%
Train Loss 0.21$\pm$0.04 0.174$\pm$0.03 0.165$\pm$0.01 0.162$\pm$0.01 0.16$\pm$0.007
Valid Loss 0.17$\pm$0.05 0.155$\pm$0.04 0.151$\pm$0.02 0.15$\pm$0.02 0.15$\pm$0.01
Table 5: KL-divergence, skewness, and kurtosis of the predictive response variable to the standard Gaussian on the ETTm2 datasets under input-96-predict-96 with $K=100$.
Variates 1 2 3 4 5 6 7
KL-Div 0.11 0.1 inf 0.13 0.03 inf 0.05
Skewness -0.42 0.41 -0.47 0.16 -0.03 -0.26 -0.72
Kurtosis 0.77 0.09 1.93 0.31 0.85 1.95 1.77
3.7 Stability of The Loss Function
We included detailed discussions and experiments on the convergence and stability of the proposed method. As shown in Table 4, the experiments show that the training and valid loss converge smoothly across multiple runs, with no evidence of instability.
Figure 8: QQ-plots and histograms of response variable.
@@ 3.8 Visualization of The Predictive Distribution
We conducted thorough statistical analyses, including QQ-plots, histograms, and metrics like KL-divergence, skewness, and kurtosis, to validate the empirical distribution of the predictive response variable. KL-divergence, skewness, and kurtosis of the predictive response variable are shown in Table 5, and QQ-plots and histograms are shown in Fig. 8. These statistical analyses shows that the empirical distribution exhibits notable characteristics, such as skewness and heavy tails, etc., which are typical of many real-world datasets.
Furthermore, the predictive distribution of DistPred is visualized in Fig. 9. It can be observed that DistPred provides an ensemble of predictions (only the top 10 are presented in the left subplot). Given all predictive ensemble values, the model can estimate the distribution of the response variable. Consequently, we can calculate confidence intervals at different levels, as shown in the right subplot of Fig. 9.
Figure 9: Visualization of the prediction results and the confidence intervals with setting input-96-predict-96 on the Ettm2 dataset. (a) The left subplot shows the ensemble of predictions with $K=100$. By utilizing subgraph (a), we can directly obtain the confidence intervals for subgraph (b), e.g, confidence intervals at 99%, 99.5% and 99.9% levels.
4 Related Work
In supervised learning contexts, the endeavor to characterize the conditional distribution $p(y|x)$ beyond merely the conditional mean $\mathop{\mathbb{E}}[y|x]$ via deep neural networks has been a focal point of existing research efforts. These endeavors primarily concentrate on quantifying uncertainty, with several approaches having been proposed.
In regression and forecasting tasks, predicting the underlying distribution of the response variable is pivotal. Traditional approaches (Bishop, 1994; Greene, 2003; Salinas et al., 2020; He et al., 2020) often rely on the assumption that the response variable conforms to a prior distribution, with the goal of estimating specific statistics derived from this distribution. Commonly, these methods transform the challenge of distribution prediction and uncertainty quantification into the prediction of statistical parameters such as the mean and variance, under the presumption of a known continuous distribution. For example, MDNs (Bishop, 1994) impose a predefined distribution—typically Gaussian—weighted by certain parameters to approximate the final distribution. Similarly, heteroscedastic regression (Greene, 2003) predicts uncertainty by modeling the variability of residuals as a function of independent variables. DeepAR (Salinas et al., 2020), another notable approach, assumes a Gaussian distribution for the response variable, thereby leveraging the GaussianNLLLoss (Nix and Weigend, 1994) to directly optimize its mean and variance. While these methods offer computational efficiency by simplifying predictions to statistical parameters, their reliance on strong distributional assumptions often limits their ability to capture the true underlying distribution, potentially leading to suboptimal performance.
Conformal prediction presents an alternative framework for distribution prediction, diverging from traditional parametric approaches. In their study, the authors in (Vovk et al., 2017, 2018) introduced a random prediction system and proposed a nonparametric prediction method grounded in conformal assumptions. By integrating conformal prediction with quantile regression in (Romano et al., 2019; Xu and Xie, 2021, 2023), they developed a method for constructing prediction intervals for the response variable. However, the practical application of conformal prediction is not without limitations. Its effectiveness is often constrained by the assumption of exchangeability of residuals, which may not hold in all contexts, particularly in the presence of temporal dependencies. This limitation can lead to less reliable prediction intervals when applied to non-independent and identically distributed (non-i.i.d.) data, thereby challenging its robustness in real-world scenarios where data often exhibit complex dependencies.
Uncertainty quantification is a method that indirectly reflects the potential distribution of the response variable. BNNs represent one such approach, aiming to capture such uncertainty by positing distributions over network parameters, thereby encapsulating the model’s plausibility given the available data (Blundell et al., 2015; Hernández-Lobato and Adams, 2015; Gal and Ghahramani, 2016; Kingma et al., 2015; Tomczak et al., 2021). Another avenue is represented by Kendall and Gal (2017), which not only addresses uncertainties in model parameters but also incorporates an additive noise term into the outputs to encompass uncertainties. In parallel, ensemble-based methodologies (Lakshminarayanan et al., 2017; Liu et al., 2022b) have emerged to address predictive uncertainty. These methods involve amalgamating multiple neural networks with stochastic outputs. Furthermore, the neural processes’ family (Garnelo et al., 2018b, a; Kim et al., 2019; Gordon et al., 2020) has introduced a suite of models tailored to capturing predictive uncertainty in a manner that extends beyond the distribution of available data, particularly tailored for few-shot learning.
@@ 5 Conclusion
In this paper, we propose a novel method named DistPred, which is a distribution-free probabilistic inference approach for regression and forecasting tasks. We transform proper scoring rules that measure the discrepancy between the predicted distribution and the target distribution into a differentiable discrete form and use it as a loss function to train the model end-to-end. This allows the model to sample numerous samples in a single forward pass to estimate the potential distribution of the response variable. Experimental results demonstrate that DistPred outperforms existing methods, often by a considerable margin. We also extend time series forecasting from point estimation to distribution prediction and achieve state-of-the-art performance on multivariate and univariate time series forecasting tasks. In the future, we plan to extend DistPred to other tasks, such as classification and reinforcement learning.
Appendix A Dataset and Implementation
A.1 Commonly Used TS Datasets
The information of the experiment datasets used in this paper are summarized as follows: (1) Electricity Transformer Temperature (ETT) dataset Zhou et al. (2021a), which contains the data collected from two electricity transformers in two separated counties in China, including the load and the oil temperature recorded every 15 minutes (ETTm) or 1 hour (ETTh) between July 2016 and July 2018. (2) Electricity (ECL) dataset  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 collects the hourly electricity consumption of 321 clients (each column) from 2012 to 2014. (3) Exchange Lai et al. (2018) records the current exchange of 8 different countries from 1990 to 2016. (4) Traffic dataset  [^2] http://pems.dot.ca.gov records the occupation rate of freeway system across State of California measured by 861 sensors. (5) Weather dataset  [^3] https://www.bgc-jena.mpg.de/wetter records every 10 minutes for 21 meteorological indicators in Germany throughout 2020. (6) Solar-Energy Lai et al. (2018) documents the solar power generation of 137 photovoltaic (PV) facilities in the year 2006, with data collected at 10-minute intervals. (7) The PEMS dataset Liu et al. (2022a) comprises publicly available traffic network data from California, collected within 5-minute intervals and encompassing 358 attributes. (8) Illness (ILI) dataset  [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html describes the influenza-like illness patients in the United States between 2002 and 2021, which records the ratio of patients seen with illness and the total number of the patients. The detailed statistics information of the datasets is shown in Table 6, and the dataset information in terms of their size and number of features is summarized in Table 7.
A.2 Implementation Details